# Energy consumption of code small language models serving with runtime engines and execution providers

### v1

@@ Abstract
Background. The rapid growth of Language Models (LMs), particularly in code generation, requires substantial computational resources, raising concerns about energy consumption and environmental impact. Optimizing LMs inference for energy efficiency is crucial, and Small Language Models (SLMs) offer a promising solution to reduce resource demands.
Aim. Our goal is to analyze the impact of deep learning runtime engines and execution providers on energy consumption, execution time, and computing-resource utilization from the point of view of software engineers conducting inference in the context of code SLMs.
Method. We conducted a technology-oriented, multi-stage experimental pipeline using twelve code generation SLMs to investigate energy consumption, execution time, and computing-resource utilization across the configurations.
Results. Significant differences emerged across configurations. CUDA execution provider configurations outperformed CPU execution provider configurations in both energy consumption and execution time. Among the configurations, TORCH paired with CUDA demonstrated the greatest energy efficiency, achieving energy savings from 37.99% up to 89.16% compared to other serving configurations. Similarly, optimized runtime engines like ONNX with the CPU execution provider achieved from 8.98% up to 72.04% energy savings within CPU-based configurations. Also, TORCH paired with CUDA exhibited efficient computing-resource utilization.
Conclusions. Serving configuration choice significantly impacts energy efficiency. While further research is needed, we recommend the above configurations best suited to software engineers’ requirements for enhancing serving efficiency in energy and performance.
@@ 1 Introduction
The widespread adoption of deep learning models, particularly Language Models (LMs), has surged remarkably in recent years. This trend has led to an exponential increase in the computational resources needed to support these models [1], raising concerns such as environmental impacts [2]. The focus within the deep learning community has predominantly been on enhancing model accuracy, often overlooking the associated costs. For instance, Benchmarks like GLUE or SuperGLUE encourage competition based on accuracy, often without much consideration for reducing energy consumption [3]. State-of-the-art models like GPT-4 and Claude 2 demand significant computational resources, primarily accessible to companies like OpenAI, Anthropic, or Google, leading to a large environmental footprint [4]. Notably, Schwartz et al. [5] emphasized the need for deep learning research to prioritize efficiency, advocating for a balance where improvements in accuracy are not achieved at the expense of a larger carbon footprint.
@@
The escalating computational demand of deep learning models, has led to an increase in research focused on improving the sustainability of these systems [6, 7]. Although the majority of this research targets the training phase, a growing body of work acknowledges the significance of both training and inference in achieving a greener deep learning [8, 9]. However, there remains a noticeable gap in studies dedicated to the inference phase[10]. This oversight is important as inference can also contribute substantially to the overall energy footprint of deep learning systems, especially when deployed at scale [11]. Addressing this gap is crucial to reduce overall environmental impact and encourage the development of energy-efficient deep learning systems.
@@
During inference, trained models are integrated into deep learning systems to allow users to make predictions, a process known as deep learning serving [12, 13]. Critical and popular architectural decisions in deep learning serving include the selection of serving infrastructure, with the runtime engine being one of the most commonly used options, and the choice of an execution provider. These decisions are driven by the community’s focus on optimizing performance and simplifying deployment to handle growing deep learning systems’ complexity [14, 15]. The runtime engine is a software component that loads the model, performs inference, and returns results to the user. It can also apply optimizations, such as graph optimizations or just-in-time compilation [16]. The execution provider, which is the library that serves as the backend of the runtime engine, optimizes model inference based on the available hardware [17, 18]. These two key decisions are able to significantly affect the energy consumption, execution time, and computing-resource utilization of the deep learning system [16].
@@
In this paper, we investigate the energy consumption of different configurations of deep learning serving infrastructures, focusing on runtime engines and execution providers. We concentrate on one particular deep learning task, specifically code generation, as it is one of the most utilized tasks of LMs and has a potential impact on software development productivity [19]. Code LMs have demonstrated impressive capabilities in code generation [20]. Their success largely stems from the Transformer architecture, which excels at learning complex patterns and scales efficiently [21, 22]. This has enabled state-of-the-art performance in models like Codex [23], the backbone of GitHub Copilot [^1] https://copilot.github.com/, a tool that provides intelligent code suggestions to developers.
@@
These advanced code LMs, however, often contain millions to billions of parameters [20, 24], requiring vast amounts of training data and substantial computational resources. In contrast, Small Language Models (SLMs) have emerged as a promising alternative, offering similar capabilities but with fewer computational demands, making them accessible to individual developers. Given their efficiency and lower resource requirements, our research focuses specifically on SLMs for code generation, a field where further exploration could lead to significant advancements [4].
@@
We evaluated the impact of various deep learning serving decisions, specifically the choice of runtime engine and execution provider, through a technology oriented experiment using twelve code SLMs. The experiment utilized a generated input dataset from the HumanEval benchmark [23], allowing us to evaluate the energy consumption, execution time, and resource utilization of each serving configuration. These configurations were based on combinations of different runtime engines and execution providers. We tested several runtime engines, including Torch, ONNX Runtime, OpenVINO Runtime, and Torch JIT, across both CPU and GPU execution providers, with the aim of identifying the most efficient serving configurations.
@@
Our results revealed notable differences in energy consumption, execution times, and resource utilization across the various serving configurations. Configurations that employed the CUDA execution provider consistently outperformed those using the CPU execution provider in both energy and time efficiency. The combination of TORCH with the CUDA execution provider, as well as ONNX or OpenVINO (OV) with the CPU execution provider, yielded the best results. TorchScript showed potential in specific use cases but still requires further refinement for more widespread application. Overall, the TORCH engine paired with the CUDA execution provider demonstrated efficient utilization of computing resources.
@@
    Actionable guidelines for software engineers: We recommend using TORCH with CUDA execution provider where GPU acceleration is feasible, and opting for optimized runtime engines such as OpenVINO runtime or ONNX runtime in CPU-only environments to achieve balanced energy and performance efficiency.
@@
    Measuring the impact of deep learning serving configurations on energy and performance: CUDA execution provider configurations consistently reduce energy consumption and execution time during inference. Using TORCH as a runtime engine and CUDA execution provider offers both optimal energy efficiency and faster execution times, with energy savings ranging from 37.99% to 89.16% and execution time reductions from 47.84% to 89.74% compared to the worst-performing runtime engine (ONNX Runtime) in CUDA execution provider configurations. The improvements are greater when evaluated against its CPU counterparts. The study emphasizes the importance of monitoring resource usage to prevent bottlenecks, such as high CPU or RAM usage in CPU execution provider configurations. It highlights the trade-offs of using optimized runtime engines, balancing energy and execution time improvements against increased resource demands, and showcases the consistent resource efficiency of the baseline runtime engine, TORCH, across various execution providers.
@@
    An analysis of deep learning serving configurations: We developed a multi-stage experimental pipeline to evaluate various deep learning serving infrastructures. This pipeline tested the energy and performance efficiency across different configurations.
Data availability statement. The replication package is publicly accessible [^2] Please refer to https://doi.org/10.5281/zenodo.14484001 following the “Cookiecutter Data Science” project structure [^3] https://cookiecutter-data-science.drivendata.org/. It includes the extracted model metadata, the developed serving infrastructure, profiling data, data analysis, statistical tests results, results for each model, and comprehensive documentation to support reproducibility and further research.
@@ 2.1 Deep learning serving configurations
In previous work [16], we identified four distinct types of deep learning serving infrastructures within the deep learning serving architectural design decisions: i) No runtime engine and web server, ii) Runtime engine and web server, iii) deep learning-specific software, and iv) End-to-end deep learning cloud service. Figure 1 illustrates the first two serving infrastructures alongside their transversal decisions. While these transversal decisions can improve the interaction between users and the ML application, they are influenced by the underlying serving infrastructure. Additionally, we include the execution provider, which acts as the backend of the runtime engine, optimizing execution for specific hardware.
Figure 1: Architectural design decisions for deep learning serving: runtime engine as a serving infrastructure and execution provider options. Figure adapted from [25].
Runtime engine. The runtime engine is a software component that loads the model into the system’s memory from a model registry, performs inference, and returns the results to the user. Optimized runtime engines (e.g. ONNX runtime, OpenVINO runtime, or Torch JIT) offer a flexible interface that facilitates the integration of hardware-specific libraries and implements model optimizations, including graph optimizations or just-in-time compilation [16, 26]. This setup can operate independently or be integrated with a web server, depending on the serving requirements. For instance, the optimized ONNX runtime can be used to load ONNX models and wrap them as a service using a web framework like FastAPI. Once the service is running, the user can make predictions through the service, using a communication protocol like gRPC (see figure 6 from [12]). Additionally, there is the option of using the deep learning framework directly without an optimized runtime engine. This is the simplest approach, as deep learning frameworks are primarily designed for training but can also be used for inference [27]. Several runtime engines are available, including:
@@
Execution providers. These are libraries that act as the back-end of the runtime engine, optimizing execution for specific hardware (CPU, GPU, FPGA, etc.) [18, 17]. There are various execution providers, including:
@@ 2.2 Code generation Small Language Models (SLMs)
Increasing evidence indicates that SLMs deliver remarkable results, especially in code-related tasks such as code generation [32, 33, 34]. This is pertinent for practitioners using limited resources or concerned with privacy issues [35, 36], situations where using Large Language Models (LLMs) is neither feasible nor desirable. Moreover, relying on closed-source LLMs like OpenAI’s ChatGPT accesible via API can introduce prohibitive costs, privacy concerns, and disproportionate energy consumption for single inferences for few tokens [37]. The exclusive capability of big tech companies to deploy state-of-the-art LLMs, due to their extensive computational and financial resources, leaves small to medium tech companies facing a significant challenge [38]. These smaller companies often operate with constrained CPU or GPU resources, necessitating innovative approaches to make use of LMs effectively. In this context, SLMs emerge as a critical solution, not only enabling a wider range of companies and deep learning practitioners to harness the power of deep learning but also contributing to environmental sustainability efforts by reducing the carbon footprint associated with large-scale computations [38].
@@ 3 Related work
There is a notable gap in studies addressing the deployment efficiency of trained deep learning models, despite inference costs potentially surpassing training due to the cumulative effect of using the model or system many times [39, 4]. Studies like Zhang et al. [40] examined hardware and software impacts on inference efficiency, while Georgiou et al. [27] compared TensorFlow and PyTorch’s energy and performance. Koubaa et al. [41] explored deployment strategies and model formats. Other works, including Hampau et al. [42] and Klimiato et al. [43], investigated containerization and deep learning serving infrastructures’ effects on energy and performance. Escribano et al. [44] focused on cloud computing energy consumption for NLP models, omitting deep learning serving infrastructure comparison.
@@
A recent exploration of the code LMs ecosystem by Yang et al. [45] offers a comprehensive analysis of code LMs, shedding light on ecosystem dynamics, model reuse strategies, and documentation practices. Their findings emphasize emerging approaches such as quantization and fine-tuning, which could directly impact deployment efficiency and robustness. In other work, Vartziotis et al. [46] compared the sustainability of the generated code itself by three LMs like GitHub Copilot, OpenAI ChatGPT-3, and Amazon CodeWhisperer. Their study emphasizes on the potential of generative LMs to contribute to environmentally conscious software development through green coding practices.
@@
Few studies have considered the runtime engine in the inference phase. Hampau [42] et al. analyzed the energy consumption from three “containerization strategies”. The difference with our paper is that we have decided to focus two different dimensions of the decisions deep learning practitioners have on the deep learning Serving Infrastructure, while they studied a single runtime engine without containerization, the same runtime engine with docker and an option without a runtime engine in a different containerization option WASM, and with a specific emphasis on Computer Vision (CV) tasks. Zhang et al. [12] conducted a large study on four different deep learning Serving Infrastructures: two deep learning-specific software and two runtime engines. Although they explored the performance of ONNX Runtime and Torch.JIT, their findings primarily reported on latency, with experiments conducted on versions containerized using Docker and communicated through gRPC API. With regard to execution providers, Alizadeh et al. [18] analyzed the performance and energy differences between CUDA and TensorRT when paired with the ONNX runtime engine and a ResNet model, finding that TensorRT outperformed CUDA. In our research, we examine the default CPU execution provider and CUDA execution provider across multiple code generation SLMs.
@@
Contrary to previous research, our study focuses on analyzing the performance of various deep learning serving configurations during the inference phase of SLM for code generation. Our analysis specifically targets runtime engines and execution providers, and their impact on code SLMs, evaluating critical factors such as energy consumption, execution time, and computing-resource utilization.

### v2

# Insights into resource utilization of code small language models serving with runtime engines and execution providers

@@ Abstract
The rapid growth of language models, particularly in code generation, requires substantial computational resources, raising concerns about energy consumption and environmental impact. Optimizing language models inference resource utilization is crucial, and Small Language Models (SLMs) offer a promising solution to reduce resource demands. Our goal is to analyze the impact of deep learning serving configurations, defined as combinations of runtime engines and execution providers, on resource utilization, in terms of energy consumption, execution time, and computing-resource utilization from the point of view of software engineers conducting inference in the context of code generation SLMs. We conducted a technology-oriented, multi-stage experimental pipeline using twelve code generation SLMs to investigate energy consumption, execution time, and computing-resource utilization across the configurations. Significant differences emerged across configurations. CUDA execution provider configurations outperformed CPU execution provider configurations in both energy consumption and execution time. Among the configurations, TORCH paired with CUDA demonstrated the greatest energy efficiency, achieving energy savings from 37.99% up to 89.16% compared to other serving configurations. Similarly, optimized runtime engines like ONNX with the CPU execution provider achieved from 8.98% up to 72.04% energy savings within CPU-based configurations. Also, TORCH paired with CUDA exhibited efficient computing-resource utilization. Serving configuration choice significantly impacts resource utilization. While further research is needed, we recommend the above configurations best suited to software engineers’ requirements for enhancing serving resource utilization efficiency.
@@ 1 Introduction
The widespread adoption of deep learning models, particularly language models, has surged remarkably in recent years. This trend has led to an exponential increase in the computational resources needed to support these models [1], raising concerns such as environmental impacts [2]. The focus within the deep learning community has predominantly been on enhancing model accuracy, often overlooking the associated costs. For instance, Benchmarks like GLUE or SuperGLUE encourage competition based on accuracy, often without much consideration for reducing energy consumption [3]. State-of-the-art models like GPT-4 and Claude 2 demand significant computational resources, primarily accessible to companies like OpenAI, Anthropic, or Google, leading to a large environmental footprint [4]. Notably, Schwartz et al. [5] emphasized the need for deep learning research to prioritize efficiency, advocating for a balance where improvements in accuracy are not achieved at the expense of a larger carbon footprint.
@@
The escalating computational demand of deep learning models, has led to an increase in research focused on improving the sustainability of these systems, including studies that evaluate deep learning architectures such as BERT, DenseNet, and RNNs in NLP and computer vision contexts [6, 7]. Although the majority of this research targets the training phase, a growing body of work acknowledges the significance of both training and inference in achieving a greener deep learning [8, 9]. However, there remains a noticeable gap in studies dedicated to the inference phase [10]. This oversight is important as inference can also contribute substantially to the overall energy footprint of deep learning systems, especially when deployed at scale [11]. Addressing this gap is crucial to reduce overall environmental impact and encourage the development of resource-efficient deep learning systems.
@@
During inference, trained models are integrated into deep learning systems to allow users to make predictions, a process known as deep learning serving [12, 13]. Critical and popular architectural decisions in deep learning serving include the selection of serving infrastructure, with the runtime engine being one of the most commonly used options, and the choice of an execution provider. These decisions are driven by the community’s focus on optimizing performance and simplifying deployment to handle growing deep learning systems’ complexity [14, 15]. The runtime engine is a software component that loads the model, performs inference, and returns results to the user. It can also apply optimizations, such as graph optimizations or just-in-time compilation [16]. The execution provider, which is the library that serves as the backend of the runtime engine, optimizes model inference based on the available hardware [17, 18]. These two key decisions are able to significantly affect the resource utilization [^1] https://din.one/site/rs. In line with the ISO 25010 [19], resource utilization is identified as a key subcharacteristic of the broader quality attribute of efficiency. This attribute refers to how effectively a system uses resources under stated conditions.
@@
In this paper, we investigate the resource utilization in terms of energy consumption, execution time, and computing-resource utilization of different configurations of deep learning serving infrastructures, focusing on runtime engines and execution providers. We concentrate on one particular deep learning task, specifically code generation, as it is one of the most popular tasks of language models and has a potential impact on software development productivity [20, 21]. As these models are increasingly integrated into development environments, optimizing their inference efficiency becomes highly relevant for practitioners concerned with both performance and sustainability. Code generation language models have demonstrated impressive capabilities in code generation [22]. Their success largely stems from the Transformer architecture, which excels at learning complex patterns and scales efficiently [23, 24]. This has enabled state-of-the-art performance in models like Codex [25], the backbone of GitHub Copilot [^2] https://copilot.github.com/, a tool that provides intelligent code suggestions to developers.
@@
These advanced code generation language models, however, often contain millions to billions of parameters [22, 26], requiring vast amounts of training data and substantial computational resources. In contrast, Small Language Models (SLMs) have emerged as a promising alternative, offering similar capabilities but with fewer computational demands, making them accessible to individual developers [27]. SLMs refer to language models with parameter counts typically below the scale of tens of billions [28, 29], a threshold that continues to evolve as the field advances [27, 30]. This distinguishes them from Large Language Models (LLMs), which often exceed tens or hundreds of billions of parameters. Code generation SLMs refer specifically to a subset of small language models trained on source code and capable of demonstrating strong performance in code generation tasks, either as their primary application or as a key evaluation benchmark. Given their efficiency and lower resource requirements, our research focuses specifically on code generation SLMs, a field where further exploration could lead to significant advancements in research and software development [4]. Specifically, we focus on the auto-complete code task, in which a SLM continuously predicts the next tokens as the developer types. This is a standard feature in modern integrated development environments (IDEs), designed to boost productivity and minimize keystrokes by generating a few tens of tokens from short input prompts [31, 32].
@@
We evaluated the impact of various deep learning serving decisions, specifically the choice of runtime engine and execution provider, through a technology-oriented experiment using twelve code generation SLMs [33]. The experiment utilized a generated input dataset from the HumanEval benchmark [25], allowing us to evaluate the energy consumption, execution time, and computing-resource utilization of each serving configuration. These configurations were based on combinations of different runtime engines and execution providers. We tested several runtime engines, including Torch (TORCH), ONNX Runtime (ONNX), OpenVINO Runtime (OV), and Torch JIT (JIT), across both CPU and GPU execution providers, with the aim of identifying the most efficient serving configurations.
@@
Our results revealed significant differences in energy consumption, execution times, and computing-resource utilization across the various serving configurations. Configurations that employed the CUDA execution provider consistently outperformed those using the CPU execution provider in both energy and time efficiency. The combination of TORCH with the CUDA execution provider, as well as ONNX or OV with the CPU execution provider, yielded the best results for those metrics. JIT can perform effectively in certain deep learning tasks, such as computer vision [34], but it needs further refinement to handle language generation scenarios such as code generation, as discussed later in this paper. Overall, the TORCH engine paired with the CUDA execution provider demonstrated efficient utilization of computing resources.
@@
    Actionable guidelines for software engineers: We draw practical guidelines for software engineers under varying serving infrastructure constraints. While these are derived from our experimental context, they are intended to support broader serving decisions in industry and academia. The guidelines address common serving trade-offs, particularly relevant to environments where sustainability, latency, or computing-resource constraints are key. These will be elaborated in detail in Section 6.
@@
    Measuring the impact of deep learning serving configurations on resource utilization: CUDA execution provider configurations consistently reduce energy consumption and execution time during inference. Using TORCH as a runtime engine and CUDA execution provider offers both optimal energy efficiency and faster execution times, with energy savings ranging from 37.99% to 89.16% and execution time reductions from 47.84% to 89.74% compared to the worst-performing runtime engine (ONNX Runtime) in CUDA execution provider configurations. The improvements are greater when evaluated against its CPU counterparts. The study emphasizes the importance of monitoring resource usage to prevent bottlenecks, such as high CPU or RAM usage in CPU execution provider configurations. It highlights the trade-offs of using optimized runtime engines, balancing energy and execution time improvements against increased resource demands, and showcases the consistent resource efficiency of the baseline runtime engine, TORCH, across various execution providers.
@@
    A reproducible pipeline and analysis of deep learning serving configurations: We developed a multi-stage experimental pipeline to evaluate various deep learning serving infrastructures. This pipeline tested the resource utilization across different configurations.
Data availability statement. The replication package is publicly accessible [^3] Please refer to https://doi.org/10.5281/zenodo.15258679 following the “Cookiecutter Data Science” project structure [^4] https://cookiecutter-data-science.drivendata.org/. It includes the extracted model metadata, the developed serving infrastructure, profiling data, data analysis, statistical tests results, results for each model, and comprehensive documentation to support reproducibility and further research.
@@ 2.1 Deep learning serving configurations
In previous work [16], we identified four distinct types of deep learning serving infrastructures within the deep learning serving architectural design decisions: i) No runtime engine and web server, ii) Runtime engine and web server, iii) deep learning-specific software, and iv) End-to-end deep learning cloud service. Figure 1 illustrates the first two serving infrastructures alongside their transversal decisions. While these transversal decisions can improve the interaction between users and the deep learning application, they are influenced by the underlying serving infrastructure. Additionally, we include the execution provider, which acts as the backend of the runtime engine, optimizing execution for specific hardware.
Figure 1: Architectural design decisions for deep learning serving: runtime engine as a serving infrastructure and execution provider options highlighted in red. Figure adapted from [35].
Runtime engine. The runtime engine is a software component that loads the model into the system’s memory from a model registry, performs inference, and returns the results to the user. Optimized runtime engines (e.g., ONNX runtime, OpenVINO runtime, or Torch JIT) offer a flexible interface that facilitates the integration of hardware-specific libraries and implements model optimizations. These include graph optimizations like constant folding and elimination of redundant nodes, as well as just in time compilation [16, 36, 37]. This setup can operate independently or be integrated with a web server, depending on the serving requirements. For instance, the optimized ONNX runtime can be used to load ONNX models and wrap them as a service using a web framework like FastAPI. Once the service is running, the user can make predictions through the service, using a communication protocol like gRPC (see Figure 6 from [12]). Additionally, there is the option of using the deep learning framework directly without an optimized runtime engine. This is the simplest approach, as deep learning frameworks are primarily designed for training but can also be used for inference [38]. Several runtime engines are available, including:
@@
Execution providers. These are libraries that act as the back-end of the runtime engine, optimizing execution for specific hardware (CPU, GPU, FPGA, etc.) [18, 17]. They perform kernel level optimizations: selecting, fusing or tuning the fine grained operations, called kernels, that span from an individual convolution to a fused block combining convolution, activation and normalization [41, 42, 43]. The specific kernel variant is chosen when the model is loaded, either through lightweight heuristics or a full exhaustive search, so that inference runs on the fastest implementation available for the detected hardware [42].
There are various execution providers, including:
@@ 2.2 Code generation Small Language Models (SLMs)
Increasing evidence indicates that SLMs deliver remarkable results in code-related tasks, such as code generation, achieving comparable accuracy and output quality to LLMs despite having significantly fewer parameters [45, 46, 47]. This is pertinent for practitioners using limited resources or concerned with privacy issues [48, 49], situations where using LLMs is neither feasible nor desirable.
In this work, we define code generation Small Language Models (SLMs) as language models that have fewer than 5 billion parameters specialized in generating source code. Our definition aligns with the perspective presented in [27], where, given specific tasks and resource constraints, SLMs fall into a range bounded by a minimum size at which the model first exhibits emergent capabilities for a specialized task, and an upper bound dictated by available computational resources. Although the exact scale of SLMs may evolve as new models emerge [30], their core objective remains stable: achieving effective performance within constrained computational conditions. Thus, in our specific context, code generation task, SLMs are models that demonstrate proficient capabilities comparable to larger models in the code generation task, but under significantly restricted computational and resource constraints compared to state-of-the-art LLMs.
Among the most prevalent real-world SLMs applications, we concentrate on the auto-complete code task of code generation, where the model observes partially written code and predicts the next few tokens, typically on the order of a few tens of tokens. This setting replicates the interactive completion facilities embedded in modern IDEs, which raise developer productivity by providing token level suggestions at the moment of typing [31]. Experience reported by practitioners further shows that conversational code assistants often deliver larger fragments, whereas auto-complete interfaces prefer more frequent but smaller suggestions. These concise few-tokens completions are easier to inspect and validate while developing, while larger blocks demand greater cognitive effort before they can be safely incorporated [31, 50].
Moreover, relying on closed-source LLMs like OpenAI’s ChatGPT accessible via API can introduce prohibitive costs, privacy concerns, and disproportionate energy consumption for single inferences for few tokens [51]. The exclusive capability of big tech companies to deploy state-of-the-art LLMs, due to their extensive computational and financial resources, leaves small to medium tech companies facing a significant challenge [52]. These smaller companies often operate with constrained CPU or GPU resources, necessitating innovative approaches to make use of language models effectively. In this context, SLMs emerge as a critical solution, not only enabling a wider range of companies and deep learning practitioners to harness the power of deep learning but also contributing to environmental sustainability efforts by reducing the carbon footprint associated with large-scale computations [52].
@@ 3 Related work
There is a notable gap in studies addressing the deployment efficiency of trained deep learning models, despite well-established evidence that inference costs can exceed training costs in deployed systems due to the cumulative effect of using the model or system many times. Reports from major technology companies such as Amazon and NVIDIA estimate that inference may account for up to 90% of total costs in production [53, 4].
@@
Studies like Zhang et al. [54] examined hardware and software impacts on inference efficiency, while Georgiou et al. [38] compared TensorFlow and PyTorch’s energy and performance. Koubaa et al. [55] explored deployment strategies and model formats. Other works, including Hampau et al. [56] and Klimiato et al. [57], investigated containerization and deep learning serving infrastructures’ effects on energy and performance. Escribano et al. [58] focused on cloud computing energy consumption for NLP models, omitting deep learning serving infrastructure comparison.
@@
A recent exploration of the code generation language models ecosystem by Yang et al. [59] offers a comprehensive analysis of code generation language models, shedding light on ecosystem dynamics, model reuse strategies, and documentation practices. Their findings emphasize emerging approaches such as quantization and fine-tuning, which could directly impact deployment efficiency and robustness.
Recent studies have examined language models for the code auto-complete task, highlighting their impact on developer productivity and system performance [31]. GitHub reports up to a 55% speed boost when using their tool GitHub Copilot in completion mode [50]. There is a growing demand to deploy these models within IDEs on local machines. However, even moderately sized models can introduce significant inference overheads like higher latency, increased energy use, and a larger carbon footprint [48]. To address these concerns, SLMs have been proposed for IDE integration [49], and dynamic inference methods, such as GREEN-CODE’s energy-aware early exits [60] and GPT-2’s first-layer predictions [61], demonstrate substantial efficiency gains when generating only a few tokens from short prompts. Yet, few works examine resource utilization during serving in the short prompt context typical of real-world auto-complete scenarios.
@@
Few studies have considered the runtime engine in the inference phase. Hampau et al. [56] analyzed the resource utilization of three different containerization strategies. Their study focused on computer vision tasks and evaluated the following strategies: a single runtime engine without containerization, the same runtime engine with Docker, and WebAssembly (WASM). In contrast, our work addresses two distinct dimensions of the deep learning serving infrastructure: the choice of runtime engine and execution provider. Zhang et al. [12] conducted a large study on four different deep learning Serving Infrastructures: two deep learning-specific software and two runtime engines. Although they explored the performance of ONNX Runtime and Torch.JIT, their findings primarily reported on latency, with experiments conducted on versions containerized using Docker and communicated through gRPC API (see Figure 1 for a reference of the different design decisions). With regard to execution providers, Alizadeh et al. [18] analyzed the performance and energy differences between CUDA and TensorRT when paired with the ONNX runtime engine and a ResNet model, finding that TensorRT outperformed CUDA. In our research, we examine the default CPU execution provider and CUDA execution provider across multiple code generation SLMs.
@@
Contrary to previous research, our study focuses on analyzing the resource utilization of various deep learning serving configurations during the inference phase of SLMs for code generation. Our analysis specifically targets runtime engines and execution providers, and their impact on code generation SLMs, evaluating critical factors such as energy consumption, execution time, and computing-resource utilization.